OpenStudy (anonymous):

expand the logarithmic expression log8 q/2 can someone explain

OpenStudy (anonymous):

\[\log(\frac{a}{b})=\log(a)-\log(b)\] is a start

OpenStudy (anonymous):

first step is \[\log_8(q)-\log_8(2)\]

OpenStudy (anonymous):

second step is to find \(\log_8(2)\) since \(\sqrt[3]{8}=2\) we have \(8^{\frac{1}{3}}\) so \(\log_8(2)=\frac{1}{3}\)

OpenStudy (anonymous):

you cannot use your calculator if you have a variable in the expression the answer is not a number, it will have a \(q\) in it